OpenStudy (anonymous):

Find all numbers for which the rational expression is undefined. -4/3x

OpenStudy (anonymous):

Only when x = 0.

OpenStudy (anonymous):

3 * 0 = 0. You cannot have a zero denominator.

OpenStudy (anonymous):

When x=0, the denominator=3*0=0, and division by zero is not defined.